OpenStudy (anonymous):

A man can clear his driveway using a snowblower in 45 minutes. It takes his son 2 hours to clear the driveway using a shovel. About how long would it take them to clear the driveway if they worked together?

OpenStudy (anonymous):

Since the man can clear his driveway in 3/4 hour, in one hour, he can clear 4/3 of his driveway. Similarly, his son can clear 1/2 of his driveway in one hour. Therefore, in one hour, working together, they can clear 4/3 + 1/2 = 11/6 of his driveway. Finally, they can clear his driveway in 6/11 hour.

OpenStudy (anonymous):

I dont really understand this!

OpenStudy (anonymous):

For this type of problem, you have to find the fraction of the job each person can do in one hour, then add them together to find the fraction (could be more than 1) of the job the can do in one hour. Finally for the entire job, the time require will be the reciprocal of the final fraction. We have the formula. Call a is the number of hours for person A, b for person B. Then the result is 1/(1/a + 1/b)

OpenStudy (shane_b):

In simplest form you just do this:\[\frac{1}{\frac{1}{45}+\frac{1}{120}}=32.73~minutes\]

OpenStudy (shane_b):

Ignore @Abs.Zero 's answer, it's incorrect :/

OpenStudy (shane_b):

BTW: This is the same as 6/11 of an hour.